# The NHIM development scenario in an asymmetric binary system of dwarf galaxies

### v1

@@ 3.4 Development scenario of the NHIM2
Finally, let us turn to NHIM2. It is created at $E_{2}\approx-0.05481$, and also this NHIM starts with an internal dynamics of a 2-dof oscillator. Also, here the horizontal Lyapunov orbit lh3 runs through a pitchfork where it splits off a pair of tilted loop orbits, it happens at $E\approx-0.046$, see again Fig.2. With increasing energy, the islands around the tilted loop orbits grow very slowly, and we must reach rather high energies to obtain some interesting plot of the restricted map of NHIM2. In Fig.5, we show the plot for E=0.035E=-0.035. Of course, since its pitchfork bifurcation, lh2 is unstable and together with the division between the central island and the tilted loop islands, it forms a separatrix structure, which, strictly speaking, is a very fine chaos strip. The apparent intersections between the various curves in the tilted loop islands are a projection effect. Here, the projection of the curved NHIM2 surface into the $y-p_{y}$ plane is no longer 1:1; however, we always want to use the same projection for all plots of the restricted map.
@@
At $E\approx-0.032$ lh2 first loses its normal hyperbolicity and immediately afterwards is destroyed in a saddle-centre bifurcation when it collides with another horizontal periodic orbit growing out of the potential minimum $P_{2}$. At $E\approx-0.02997$, the tilted loop orbits in NHIM2 first lose normal hyperbolicity and immediately afterwards are destroyed in saddle-centre bifurcations when they collide with other tilted loop orbits grown out of $P_{2}$, see again Fig.2. The vertical orbit lv2 remains normally hyperbolic up to very high energy, and accordingly, also a remnant of the NHIM2 around lv2 remains. Qualitatively, the scenario of NHIM2 is a mirror image of the scenario of NHIM3, even though the analogous individual events happen at higher values of $E$. So we see in which sense a qualitative mirror symmetry of the system is conserved, also when the symmetry between the masses is broken.
3.5 Some comments on the NHIM development scenarios
@@
Important events happen in NHIM1 and also in NHIM3 in the same relatively small energy interval between E=0.043E=-0.043 and E=0.04E=-0.04. With increasing energy in NHIM3, the loss of normal hyperbolicity happens before the development of large-scale chaos in the internal dynamics. We see large-scale chaos in the surroundings of the remnant parts of NHIM3 only as a transient effect. In contrast, in the internal dynamics of NHIM1, we observe the development of large-scale chaos before the beginning of any decay of this NHIM. The difference is that NHIM3 starts its decay because the normal instability is becoming small, whereas in NHIM1, the development of the growing tangential instability is important for the beginning of the decay.
@@
When we compare the development scenario of NHIM1 on one hand with the development scenarios of NHIM2 and NHIM3 on the other hand, then we notice the following qualitative difference. On NHIM1, the horizontal Lyapunov orbit runs through two consecutive pitchfork bifurcations, where in the second one, a pair of unstable tilted loop orbits is created, while the horizontal Lyapunov orbit returns to tangential stability. This second pair of tilted loop orbits turns vertical and is finally absorbed by the vertical Lyapunov orbit in an inverse pitchfork bifurcation. Thereby, the vertical Lyapunov orbit becomes tangentially unstable, and this event triggers the creation of large-scale chaos inside of NHIM1.
@@
In contrast, in NHIM2 and in NHIM3, tangentially unstable tilted loop orbits are never created, which has two consequences. First, the horizontal Lyapunov orbit remains tangentially unstable, which favours the beginning of the decay of these NHIMs from the separatrix structure, which contains lh2 or lh3, respectively. Second, the vertical Lyapunov orbit remains tangentially stable, which favours the survival of the central island of these NHIMs up to very high energy.
@@
We made a surprising observation of the connection of these properties with the behaviour of the time periods of the Lyapunov orbits as a function of the energy. In Fig.6, we plot these time periods as a function of the energy. As before, the red curves belong to the horizontal Lyapunov orbits, and the green curves belong to the vertical Lyapunov orbits. The blue curves belong to the tangentially stable tilted loop orbits. And the magenta curve belongs to the tangentially unstable tilted loop orbits. The solid curves belong to NHIM1, the dotted curves belong to NHIM2 and the broken curves belong to NHIM3. In NHIM1, the periods of the horizontal and the vertical Lyapunov orbit cross each other; one pitchfork bifurcation happens energetically before the crossing, and the other one happens after the crossing. In NHIM2 and NHIM3, we do not find any analogous crossing of the horizontal and vertical periods.
@@
Figure 6: Periods of Lyapunov orbits and tilted loop orbits as a function of the energy.
@@
Now the reader might ask whether this observation is a strange feature of the particular example or whether it is common. Interestingly, the experience with other systems suggests that this behaviour is common. In the system of a test particle in a barred galaxy [27] and for the dynamics of a test particle in a double-barred galaxy [22], we found two consecutive pitchfork bifurcations creating first a pair of tangentially stable tilted loop orbits and later a pair of tangentially unstable tilted loop orbits. Also, a crossing of the periods of the horizontal and the vertical Lyapunov orbits occurs between these two pitchfork bifurcations. However, in these two examples, the tilted loop orbits come out of the vertical Lyapunov orbit. On the other hand, in the system of a test particle in the gravitational field of an open star cluster [28], there is no crossing of periods, and there is only a single pitchfork bifurcation creating a pair of tangentially stable tilted loop orbits.
@@
In a classical model for a hydrogen atom in a rotating external field [29], we find the following scenario. First the horizontal Lyapunov orbit splits off a pair of tangentially stable tilted loop orbits. Later, we have a crossing of the periods of the horizontal and vertical Lyapunov orbits. However, at an energy a little higher, the horizontal Lyapunov orbit collides with the singularity of the Coulomb system at the origin, and this event throws the scenario out of its usual path. At an energy still a little higher, the horizontal Lyapunov orbit is absorbed by an orbit coming out of the central potential hole in an inverse period doubling bifurcation. So the horizontal Lyapunov orbit ends its existence before it has the opportunity for a further pitchfork bifurcation.
@@
The conclusion of this subsection is the following. We have seen the creation of tangentially unstable tilted loop orbits only in cases where the periods of the horizontal and the vertical Lyapunov orbits cross. The system of the two dwarf galaxies is a nice example where we see in the same system the case with crossing and the case without crossing. Certainly, the crossing of the periods indicates some kind of 1:1 resonance between the horizontal and the vertical motion. And such a resonance is needed for the creation of tilted loop orbits out of the Lyapunov orbits. Unfortunately, so far we are not yet able to give a detailed explanation of all these connections.
@@ 4 Heteroclinic connections and the appearance of large-scale chaos
Now we relate the NHIM development scenario seen in section 3 to some particularly interesting classes of trajectories and their properties, namely, to heteroclinic trajectories between the 3 NHIMs. We do it for the energy values 0.043-0.043, 0.042-0.042 and 0.04-0.04, i.e. exactly for the energy range where the most interesting events of the whole development scenario happen. Because of the easier pictorial presentation, we begin with horizontal trajectories. The only horizontal trajectories within the NHIMs are the respective horizontal Lyapunov orbits, which have already served as boundaries of the NHIM representations in the plots of the restricted map. Fig.7 shows the three horizontal Lyapunov orbits lh1, lh2 and lh3 (plotted as red lines) in the horizontal plane of the position space for E=0.043E=-0.043. These three orbits are identified in the plot by their order along the $x$ axis. We know already from Fig.2 that this energy value is very close to the highest energy where all the important fundamental periodic orbits still exist. The figure contains, in addition, the projection into the horizontal plane of the vertical Lyapunov orbits (dark green curves) and of the tilted loop orbits (blue curves). The orbit lu1 disappears already exactly at this energy when it collides with lv1. Therefore, it looks very similar to lv1 at this energy, and it is not included in the figure because it would be extremely difficult to distinguish lu1 and lv1. Note that lh3 and lh1 come very close to each other, while lh2 keeps a greater distance from lh1. Note that the two tilted loop orbits of each NHIM have identical projections into the horizontal plane.
Figure 7: Projection of the simplest and most important periodic orbits at E=0.043E=-0.043 into the horizontal plane of the position space. The red curves are horizontal Lyapunov orbits, the green curves are vertical Lyapunov orbits , and the blue curves are tangentially stable tilted loop orbits.
@@
4.1 Heteroclinic trajectories
@@
Figure 8: Heteroclinic trajectories between various substructures of the NHIMs at E=0.043E=-0.043. (a) : A horizontal heteroclinic trajectory going from lh3 to lh1. (b) : A horizontal heteroclinic trajectory going from lh3 to lh2. (c) : A horizontal heteroclinic trajectory going from lh2 to lh1. (d) : Projection into the horizontal plane of a heteroclinic trajectory going from lv3 to the outer part of the tilted loop island of NHIM1. (e) : Projection into the horizontal plane of a heteroclinic trajectory going from lt3 to the outer part of the central island of NHIM2.
@@
First, we have a look at heteroclinic orbits between the horizontal Lyapunov orbits. They are the only horizontal heteroclinic orbits between the various NHIMs. In Fig.8(a), we show a typical example of a heteroclinic trajectory which starts on lh3 and ends on lh1. In Fig.8(b), we show a heteroclinic trajectory going from lh3 to lh2 and Fig.8(c) presents a heteroclinic trajectory going from lh2 to lh1. Of course, there is an infinity of further horizontal heteroclinic trajectories between the various horizontal Lyapunov orbits performing all kinds of loops in the regions of the NHIMs before settling down to one particular Lyapunov orbit in the future and in the past.
@@
We also show two examples of heteroclinic trajectories going out of the horizontal plane. However, in the figures, we plot their projections into the horizontal plane. Fig.8(d) presents a trajectory starting on lv3 and ending on the outer part of the tilted loop island of NHIM1 close to the separatrix. Fig. 8(e) presents a trajectory starting on lt3 and ending on the outer part of the central island of NHIM2.
@@
Now we increase the energy a little to -0.042. We know already that for this energy, the periodic orbit lh3 has already disappeared in a saddle-centre bifurcation included in Fig.2. Accordingly, there is no longer any horizontal trajectory inside the remaining part of NHIM3 and therefore also no horizontal heteroclinic trajectory starting or ending on NHIM3. However, we have the transient chaotic region presented in the NHIM3 plot of Fig.4(c). And some of these transient trajectories around the remnants of NHIM3 have a projection into the horizontal plane which qualitatively looks similar to lh3 at smaller energies, i.e. it looks as if there would be a logical continuation of lh3 to higher energies. From the neighbourhood of such transient trajectories, heteroclinic trajectories to other NHIMs can start. Fig.9(a) shows the projection into the horizontal plane of the position space of a typical trajectory going from the transient part of NHIM3 to NHIM2. Its initial point lies very close to the 26th point of the transient trajectory shown in Fig.4(c), its coordinates in this figure are $y\approx 0.57,p_{y}\approx-0.39$. On the NHIM2, this heteroclinic trajectory approaches the outer part of the central island in NHIM2.
@@
Fig.9(b) shows the projection into the horizontal plane of the position space of a trajectory going from the transient part of NHIM3 to NHIM1. The initial point of this trajectory lies extremely close to the initial point of the trajectory shown in Fig.9(a). Compare in the figures the initial point and the first loop of these two trajectories in the region of NHIM3. However, in the long run, these two trajectories are completely different. This demonstrates the extreme dependence on the initial conditions of trajectories starting in the chaotic region, a typical indication of chaos. This heteroclinic trajectory approaches the NHIM1 in the region between the moderate-sized chaos strip created around the stable and unstable manifolds of the central hyperbolic fixed point corresponding to the vertical Lyapunov orbit lv1 and the outermost layer of the tilted loop islands, which is already broken into secondary structures and chaos strips. It establishes a connection between the transient chaotic and decaying part of NHIM3 and the part of NHIM1 which is just in the process of turning chaotic at this stage of the development scenario.
@@
Fig.9(c) shows a trajectory starting in the chaos region of NHIM1 and coming close to the outer transient parts of NHIM3. Fig.9(d) shows a trajectory starting close to lv1, i.e. in the centre of the chaotic region of NHIM1 and approaching NHIM2.
@@
Figure 9: Projection into the horizontal plane of four heteroclinic trajectories between the NHIMs at E=0.042E=-0.042. Part (a): A heteroclinic trajectory from the transient part of NHIM3 to NHIM2. part (b): A heteroclinic trajectory from the transient part of NHIM3 to NHIM1. part (c): A heteroclinic trajectory from the large-scale chaotic region of NHIM1 to NHIM3. part (d): A heteroclinic trajectory from lv1 to NHIM2.
@@
Next, we go to E=0.04E=-0.04 where NHIM3 and NHIM1 have both started with their decay. The Fig.10(a) shows a heteroclinic trajectory starting on the transient part of NHIM3 and going to NHIM1. And the Fig.10(b) shows a heteroclinic trajectory starting on the transient part of NHIM1 and ending on NHIM3.
@@
Figure 10: Projection into the horizontal plane of two trajectories at E=0.04E=-0.04. Part (a): A heteroclinic trajectory going from the transient part of NHIM3 to NHIM1. Part (b): A homoclinic trajectory going from the transient part of NHIM1 to NHIM3.

### v2

@@ 3.4 Bifurcation scenario of the NHIM $\mathcal{M}^{2}_{E_{J}}$
Finally, let us turn to NHIM $\mathcal{M}^{2}_{E_{J}}$. It is created at $E_{J2}\approx-0.05481$, and also this NHIM starts with the internal dynamics of a 2-dof oscillator. Also, here the horizontal Lyapunov orbit $lh_{2}$ runs through a pitchfork where it splits off a pair of tilted loop orbits, it happens at $E_{J}\approx-0.046$, see Fig.2. With increasing Jacobi constant, the KAM tori around the tilted loop orbits grow very slowly, and we must reach rather high energies to obtain some interesting plot of the projected Poincaré map of NHIM $\mathcal{M}^{2}_{E_{J}}$. In Fig.5, we show the plot for $E_{J}=-0.035$. Of course, since its pitchfork bifurcation, $lh_{2}$ is unstable, and together with the division between the central KAM island and the tilted loop islands in the Poincaré map, it forms a separatrix structure, which, strictly speaking, is a very fine chaos strip. The apparent intersections between the various curves in the tilted loop islands are a projection effect. Here, the projection of the curved NHIM $\mathcal{M}^{2}_{E_{J}}$ surface into the $yp_{y}$ plane is no longer 1:1; however, we always want to use the same projection for all plots.
@@
At $E_{J}\approx-0.032$ $lh_{2}$ first loses its normal hyperbolicity and immediately afterwards is destroyed in a saddle-centre bifurcation when it collides with another horizontal periodic orbit growing out of the potential minimum $P_{2}$. At $E_{J}\approx-0.02997$, the tilted loop orbits in NHIM $\mathcal{M}^{2}_{E_{J}}$ first lose normal hyperbolicity and immediately afterwards are destroyed in saddle-centre bifurcations when they collide with other tilted loop orbits grown out of $P_{2}$, see Fig.2. The vertical orbit $lv_{2}$ remains normally hyperbolic up to very high Jacobi constant, and accordingly, also a remnant of the NHIM $\mathcal{M}^{2}_{E_{J}}$ around $lv_{2}$ remains. Qualitatively, the scenario of NHIM $\mathcal{M}^{2}_{E_{J}}$ is a mirror image of the scenario of NHIM $\mathcal{M}^{3}_{E_{J}}$, even though the analogous individual events happen at higher values of $E_{J}$. So we see in which sense a qualitative mirror symmetry of the system is conserved, also when the symmetry between the masses is broken.
3.5 Some comments on the NHIM bifurcation scenarios
@@
Important events happen in NHIM $\mathcal{M}^{1}_{E_{J}}$ and also in NHIM $\mathcal{M}^{3}_{E_{J}}$ in the same relatively small Jacobi constant interval between E=0.043E=-0.043 and $E_{J}=-0.04$. With increasing Jacobi constant in NHIM $\mathcal{M}^{3}_{E_{J}}$, the loss of normal hyperbolicity happens before the development of large-scale chaos in the internal dynamics. We see large-scale chaos in the surroundings of the remnant parts of NHIM $\mathcal{M}^{3}_{E_{J}}$ only as a transient effect. In contrast, in the internal dynamics of NHIM $\mathcal{M}^{1}_{E_{J}}$, we observe the development of large-scale chaos before the beginning of any break of this NHIM. The difference is that NHIM $\mathcal{M}^{3}_{E_{J}}$ starts its break because the normal instability is becoming small, whereas in NHIM $\mathcal{M}^{1}_{E_{J}}$, the growth of the tangential instability is important for the beginning of the break.
@@
When we compare the bifurcation scenario of NHIM $\mathcal{M}^{1}_{E_{J}}$ on one hand with the bifurcation scenarios of NHIM $\mathcal{M}^{2}_{E_{J}}$ and NHIM $\mathcal{M}^{3}_{E_{J}}$ on the other hand, then we notice the following qualitative difference. On NHIM $\mathcal{M}^{1}_{E_{J}}$, the horizontal Lyapunov orbit runs through two consecutive pitchfork bifurcations, where in the second one, a pair of unstable tilted loop orbits is created, while the horizontal Lyapunov orbit returns to tangential stability. This second pair of tilted loop orbits turns vertical and is finally absorbed by the vertical Lyapunov orbit in an inverse pitchfork bifurcation. Thereby, the vertical Lyapunov orbit becomes tangentially unstable, and this event triggers the creation of large-scale chaos inside of NHIM $\mathcal{M}^{1}_{E_{J}}$.
@@
In contrast, in NHIM $\mathcal{M}^{2}_{E_{J}}$ and in NHIM $\mathcal{M}^{3}_{E_{J}}$, tangentially unstable tilted loop orbits are never created, which has two consequences. First, the horizontal Lyapunov orbit remains tangentially unstable, which favours the beginning of the break of these NHIMs from the separatrix structure, which contains $lh_{2}$ or $lh_{3}$, respectively. Second, the vertical Lyapunov orbit remains tangentially stable, which favours the survival of the central KAM island of the Poincaré map of these NHIMs up to very high Jacobi constant.
@@
We made a surprising observation of the connection between these properties and the behaviour of the time periods of the Lyapunov orbits as a function of the Jacobi constant. In Fig.6, we plot these time periods as a function of the Jacobi constant. As before, the red curves belong to the horizontal Lyapunov orbits, and the green curves belong to the vertical Lyapunov orbits. The blue curves belong to the tangentially stable tilted loop orbits. And the magenta curve belongs to the tangentially unstable tilted loop orbits. The solid curves belong to NHIM $\mathcal{M}^{1}_{E_{J}}$, the dotted curves belong to NHIM $\mathcal{M}^{2}_{E_{J}}$ and the broken curves belong to NHIM $\mathcal{M}^{3}_{E_{J}}$. In NHIM $\mathcal{M}^{1}_{E_{J}}$, the periods of the horizontal and the vertical Lyapunov orbit cross each other; one pitchfork bifurcation happens energetically before the crossing, and the other one happens after the crossing. In NHIM $\mathcal{M}^{2}_{E_{J}}$ and NHIM $\mathcal{M}^{3}_{E_{J}}$, we do not find any analogous crossing of the horizontal and vertical periods.
@@
Figure 6: Periods of Lyapunov orbits and tilted loop orbits as a function of the Jacobi constant $E_{J}$.
@@
Now the reader might ask whether this observation is a strange feature of the particular example or whether it is common. Interestingly, the experience with other systems suggests that this behaviour is common. In the system of a test particle in a barred galaxy [29] and for the dynamics of a test particle in a double-barred galaxy [24], there are two consecutive pitchfork bifurcations creating first a pair of tangentially stable tilted loop orbits and later a pair of tangentially unstable tilted loop orbits. Also, a crossing of the periods of the horizontal and the vertical Lyapunov orbits occurs between these two pitchfork bifurcations. However, in these two examples, the tilted loop orbits come out of the vertical Lyapunov orbit. On the other hand, in the system of a test particle in the gravitational field of an open star cluster [30], there is no crossing of periods, and there is only a single pitchfork bifurcation creating a pair of tangentially stable tilted loop orbits.
@@
In a classical model for a hydrogen atom in a rotating external field [31], the scenario is as follows. First the horizontal Lyapunov orbit splits off a pair of tangentially stable tilted loop orbits. Later, there is a crossing of the periods of the horizontal and vertical Lyapunov orbits. However, at a Jacobi constant a little higher, the horizontal Lyapunov orbit collides with the singularity of the Coulomb system at the origin, and this event throws the scenario out of its usual path. At a Jacobi constant still a little higher, the horizontal Lyapunov orbit is absorbed by an orbit coming out of the central potential hole in an inverse period doubling bifurcation. So the horizontal Lyapunov orbit ends its existence before it has the opportunity for a further pitchfork bifurcation.
@@
The conclusion of this subsection is the following. We have seen the creation of tangentially unstable tilted loop orbits only in cases where the periods of the horizontal and the vertical Lyapunov orbits cross. The system of the two dwarf galaxies is a nice example where we see in the same system the case with crossing and the case without crossing. Certainly, the crossing of the periods indicates some kind of 1:1 resonance between the horizontal and the vertical motion. And such a resonance is needed for the creation of tilted loop orbits out of the Lyapunov orbits. Unfortunately, so far, we are not yet able to give a detailed explanation of all these connections.
@@ 4 Heteroclinic connections and the appearance of large-scale chaos
Now we relate the NHIM bifurcation scenarios seen in section 3 to some particularly interesting classes of trajectories and their properties, namely, to heteroclinic trajectories between the 3 NHIMs.We do it for the Jacobi constant values $E_{J}=-0.043$, 0.042-0.042 and 0.04-0.04, i.e. exactly for the Jacobi constant range where the most interesting events of the whole bifurcation scenario happen. Because of the easier pictorial presentation, we begin with horizontal trajectories. The only horizontal trajectories within the NHIMs are the respective horizontal Lyapunov orbits, which have already served as boundaries of the NHIM representations in the plots of the projected Poincaré map. Fig.7 shows the three horizontal Lyapunov orbits $lh_{1}$, $lh_{2}$ and $lh_{3}$ (plotted as red lines) in the horizontal plane of the position space for E=0.043E=-0.043. These three orbits are identified in the plot by their order along the $x$ axis. We know already from Fig.2 that this Jacobi constant value is very close to the highest Jacobi constant where all the important fundamental periodic orbits still exist. The figure contains, in addition, the projection into the horizontal plane of the vertical Lyapunov orbits (dark green curves) and of the tilted loop orbits (blue curves). The orbit $lu_{1}$ disappears already exactly at this Jacobi constant when it collides with $lv_{1}$. Therefore, it looks very similar to $lv_{1}$ at this Jacobi constant, and it is not included in the figure because it would be extremely difficult to distinguish $lu_{1}$ and $lv_{1}$. Note that $lh_{3}$ and $lh_{1}$ come very close to each other, while $lh_{2}$ keeps a greater distance from $lh_{1}$. The two tilted loop orbits of each NHIM have identical projections into the horizontal plane.
Figure 7: Projection of the simplest and most important periodic orbits at $E_{J}=-0.043$ into the horizontal plane of the position space. The red curves are horizontal Lyapunov orbits, the green curves are vertical Lyapunov orbits, and the blue curves are tangentially stable tilted loop orbits.
@@
4.1 Heteroclinic trajectories and transient dynamics associated with the NHIMs break
@@
Figure 8: Heteroclinic trajectories between various substructures of the NHIMs at $E_{J}=-0.043$. (a) : A horizontal heteroclinic trajectory going from $lh_{3}$ to $lh_{1}$. (b) : A horizontal heteroclinic trajectory going from $lh_{3}$ to $lh_{2}$. (c) : A horizontal heteroclinic trajectory going from $lh_{2}$ to $lh_{1}$. (d) : Projection into the horizontal plane of a heteroclinic trajectory going from $lv_{3}$ to the outer part of the tilted loop island of NHIM $\mathcal{M}^{1}_{E_{J}}$. (e) : Projection into the horizontal plane of a heteroclinic trajectory going from $lt_{3}$ to the outer part of the central island of NHIM $\mathcal{M}^{2}_{E_{J}}$.
@@
First, we have a look at heteroclinic orbits between the horizontal Lyapunov orbits. They are the only horizontal heteroclinic orbits between the various NHIMs. In Fig.8(a), we show a typical example of a heteroclinic trajectory which starts on $lh_{3}$ and ends on $lh_{1}$. In Fig.8(b), we show a heteroclinic trajectory going from $lh_{3}$ to $lh_{2}$ and Fig.8(c) presents a heteroclinic trajectory going from $lh_{2}$ to $lh_{1}$. Of course, there is an infinity of further horizontal heteroclinic trajectories between the various horizontal Lyapunov orbits perform all kinds of loops in the regions of the NHIMs before settling down to one particular Lyapunov orbit in the past and in the future.
@@
We also show two examples of heteroclinic trajectories going out of the horizontal plane. However, in the figures, we plot their projections into the horizontal plane. Fig.8(d) presents a trajectory starting on $lv_{3}$ and ending on the outer part of the tilted loop island of NHIM $\mathcal{M}^{1}_{E_{J}}$ close to the separatrix. Fig. 8(e) presents a trajectory starting on $lt_{3}$ and ending on the outer part of the central island of NHIM $\mathcal{M}^{2}_{E_{J}}$.
@@
Now we increase the Jacobi constant a little to $E_{J}=-0.042$. We know already that for this Jacobi constant, the periodic orbit $lh_{3}$ has already disappeared in a saddle-centre bifurcation included in Fig.2. Accordingly, there is no longer any horizontal trajectory inside the NHIM $\mathcal{M}^{3}_{E_{J}}$ and therefore also no horizontal heteroclinic trajectory starting or ending on NHIM $\mathcal{M}^{3}_{E_{J}}$. However, we have the transient chaotic region close to the NHIM $\mathcal{M}^{3}_{E_{J}}$ on the plot of Fig.4(c). And some of these transient trajectories close to NHIM $\mathcal{M}^{3}_{E_{J}}$ have a projection into the horizontal plane which qualitatively looks similar to $lh_{3}$ at smaller energies, i.e. it looks as if there would be a logical continuation of $lh_{3}$ to higher energies. From the neighbourhood of such transient trajectories, other trajectories can go to other NHIMs. Fig.9(a) shows the projection into the horizontal plane of the position space of a typical trajectory going from the transient region close to NHIM $\mathcal{M}^{3}_{E_{J}}$ to NHIM $\mathcal{M}^{2}_{E_{J}}$. Its initial point lies very close to the 26th Poincaré map iteration of the transient trajectory shown in Fig.4(c), its coordinates in this figure are $y\approx 0.57,p_{y}\approx-0.39$. On the NHIM $\mathcal{M}^{2}_{E_{J}}$, this transient chaotic trajectory approaches the outer part of the central island in NHIM $\mathcal{M}^{2}_{E_{J}}$.
@@
Fig.9(b) shows the projection into the horizontal plane of the position space of a trajectory going from the transient region close to NHIM $\mathcal{M}^{3}_{E_{J}}$ to NHIM $\mathcal{M}^{1}_{E_{J}}$. The initial point of this trajectory lies extremely close to the initial point of the trajectory shown in Fig.9(a). Compare in the figures the initial point and the first loop of these two trajectories in the region of NHIM $\mathcal{M}^{3}_{E_{J}}$. However, in the long run, these two trajectories are completely different. This demonstrates the extreme dependence on the initial conditions of trajectories starting in the chaotic region, a typical indication of chaos. This trajectory approaches the NHIM $\mathcal{M}^{1}_{E_{J}}$ in the region between the moderate-sized chaos strip created around the stable and unstable manifolds of the central hyperbolic fixed point corresponding to the vertical Lyapunov orbit $lv_{1}$ and the outermost layer of the tilted loop islands, which is already broken into secondary structures and chaos strips. It establishes a connection between the transient chaotic region around NHIM $\mathcal{M}^{3}_{E_{J}}$ and the part of NHIM $\mathcal{M}^{1}_{E_{J}}$, which is just in the process of turning chaotic at this stage of the bifurcation scenario.
@@
Fig.9(c) shows a trajectory starting in the chaotic region of NHIM $\mathcal{M}^{1}_{E_{J}}$ and coming close to the transient chaotic region close to NHIM $\mathcal{M}^{3}_{E_{J}}$. Fig.9(d) shows a trajectory starting close to $lv_{1}$, i.e. in the centre of the chaotic region of NHIM $\mathcal{M}^{1}_{E_{J}}$ and approaching NHIM $\mathcal{M}^{2}_{E_{J}}$.
@@
Figure 9: Projection into the horizontal plane of four trajectories between the neighbourhoods of NHIMs at $E_{J}=-0.042$. Part (a): A trajectory from the transient part close to NHIM $\mathcal{M}^{3}_{E_{J}}$ to NHIM $\mathcal{M}^{2}_{E_{J}}$. part (b): A trajectory from the transient part close to NHIM $\mathcal{M}^{3}_{E_{J}}$ to NHIM $\mathcal{M}^{1}_{E_{J}}$. part (c): A trajectory from the large-scale chaotic region of NHIM $\mathcal{M}^{1}_{E_{J}}$ to NHIM $\mathcal{M}^{3}_{E_{J}}$. part (d): A heteroclinic from the central region of NHIM $\mathcal{M}^{1}_{E_{J}}$, close to the trajectory from $lv_{1}$, to NHIM $\mathcal{M}^{2}_{E_{J}}$.
@@
Next, let us consider $E_{J}=-0.04$ where NHIM $\mathcal{M}^{3}_{E_{J}}$ and NHIM $\mathcal{M}^{1}_{E_{J}}$ have both started with their break. The Fig.10(a) shows a trajectory starting in the transient region close to NHIM $\mathcal{M}^{3}_{E_{J}}$ and going to NHIM $\mathcal{M}^{1}_{E_{J}}$. And the Fig.10(b) shows a trajectory starting on the transient part of NHIM $\mathcal{M}^{1}_{E_{J}}$ and ending on NHIM $\mathcal{M}^{3}_{E_{J}}$.
@@
Figure 10: Projection into the horizontal plane of two trajectories at $E_{J}=-0.04$. Part (a): A trajectory going from the transient region close to NHIM $\mathcal{M}^{3}_{E_{J}}$ to NHIM $\mathcal{M}^{1}_{E_{J}}$. Part (b): A trajectory going from the transient region close to NHIM $\mathcal{M}^{1}_{E_{J}}$ to NHIM $\mathcal{M}^{3}_{E_{J}}$.